OpenStudy (anonymous):

Choose the equation below that represents the line that passes through the point (2, 4) and has a slope of 3. y - 4 = 3(x - 2) y - 2 = 3(x - 4) y + 4 = 3(x + 2) y + 2 = 3(x + 4)

OpenStudy (anonymous):

Okay so for this one, all you have to do is look at where the numbers are placed. Here is the equation for this type of problem. y-y(1) = m(x-x(1)) Where m= slope, which is 3. And y(1)= 4, and x(1)= 2... Do you get it?

OpenStudy (anonymous):

umm... haha no sorry

OpenStudy (anonymous):

Its all good! What I am trying to say is this is an equation that is already filled out for you. They put everything in its place. So given the values I gave you, and the original equation od you think you could fill it out?

OpenStudy (anonymous):

ahhh!! I suck at math:/

OpenStudy (anonymous):

 Thats how I would do it... Just plug in the values into the equation.